Question 1 22 - 90° A 220 V 1290° V j32 322 -j52 Find the SAEs for the assigned loop currents. Solve for the loop currents.
Answer #1

Page e 13 24-909A. dolna 2lov TO 1 Loov jer & 12 SAR JOT isr for Loop (nots in Loop 2, Ia cunant files I = 22-90° AT in hoofum Qov sun 2 2980 at node Nox 220° v.) at node 1 VL-V3 V-2 2/90°=0. -js 2V-2V3 +63(2-2) = 20 V7(2-js) -2 V3 = 20 – 10j o at n-T 148-U X (18) + Too = 12 V1 Vt = 3-88/34-39E